OpenStudy (anonymous):

Isolate y in the equation 3x + 2y = 8.

OpenStudy (anonymous):

isolating y in the equation just means manipulate the equation into the form y = (8 - 3x)/2 @quinjenkins

OpenStudy (alesia/bailey):

3x+2y=8 -3x 2y=8-3x /2 y=4-3x/2